The relation between time t and distance x is t=ax^2+ BC where a and B are constants. Find the acceleration in terms of velocity(v).

Grade:11

1 Answers

Himanshu Rathour
5 years ago
Given
t=ax²+bc
Differentiating above equation wrt x
1=2ax(dx/dt)
dx/dt=v
1/2ax=v.....i
Differentiating above equation wrt x
dv/dx=-1/ax².....ii
Now
A=dv/dt
Multiplying and dividing by dx
A=v(dv/dx) .....iii
From i ii & iii
A=-1/2ax3